# Restore-RWKV: Efficient and Effective Medical Image Restoration with RWKV

### v2

@@ 4.3 Evaluation
Three popular metrics are used for quantitative comparisons: Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM), and Root Mean Squared Error (RMSE). Higher values of PSNR and SSIM indicate better performance, while lower values of RMSE signify better results. Additionally, we provide comparisons of model size and FLOPs. The input image size is set to 128$\times$128 when calculating FLOPs.
@@ 5 Experimental Results
5.1 MRI Image Super-Resolution Results
For MRI image super-resolution, we compare our proposed Restore-RWKV with three specific image super-resolution methods: SRCNN [1], VDSR [2], and EDSR [3], as well as four general image restoration methods: SwinIR [28], Restormer [43], Xformer [44], and MambaIR [29]. Table 1 demonstrates that Restore-RWKV achieves the best performance in MRI image super-resolution while maintaining good efficiency. Restore-RWKV attains an effective global receptive field with linear computational complexity, thereby better balancing computational efficiency and model performance compared to all CNN-based, Transformer-based, and Mamba-based methods presented in Table 1. Notably, Restore-RWKV outperforms the second-best method Restormer, by more than 0.20 dB in PSNR. Figure 5 (a) illustrates visualization results across different methods, showing that our Restore-RWKV excels in recovering structures and details in MRI images.
@@ 5.1 MRI Image Super-Resolution Results
Method Type Params (M) FLOPs (G) PSNR↑ SSIM↑ RMSE↓
SRCNN CNN 0.0573 0.9369 28.8067 0.8919 41.3488
VDSR CNN 0.6647 10.8905 30.0446 0.9140 36.0508
EDSR CNN 38.3432 628.2145 31.4168 0.9338 31.1390
SwinIR Transformer 11.4977 187.9274 31.5549 0.9334 30.5788
@@
Xformer Transformer 25.2197 35.7291 31.5697 0.9338 30.5616
MambaIR Mamba 31.5035 34.3491 31.7677 0.9369 29.8372
Restore-RWKV RWKV 27.9140 37.4585 32.0913 0.9408 28.9713
[…]
Method Type Params (M) FLOPs (G) PSNR↑ SSIM↑ RMSE↓
REDCNN CNN 1.8489 24.9226 33.1889 0.9113 8.9427
Eformer Transformer 0.3441 1.2775 33.3496 0.9175 8.8030
CTformer Transformer 1.4475 3.4437 33.2506 0.9134 8.8974
SwinIR Transformer 11.4977 187.9274 33.6844 0.9179 8.4650
Restormer Transformer 26.1241 35.2052 33.6610 0.9182 8.4909
Xformer Transformer 25.2197 35.7291 33.6903 0.9183 8.4621
[…]
Restore-RWKV RWKV 27.9140 37.4585 33.7988 0.9198 8.3600
[…]
Method Type Params (M) FLOPs (G) PSNR↑ SSIM↑ RMSE↓
Xiang’s CNN 0.2293 3.7654 35.9268 0.9167 0.0980
DCNN CNN 1.3331 21.8576 36.2710 0.9243 0.0954
ARGAN CNN 31.1426 8.4807 36.7272 0.9406 0.0902
SwinIR Transformer 11.4977 187.9274 37.1559 0.9453 0.0868
Restormer Transformer 26.1241 35.2052 37.1617 0.9460 0.0869
@@
Restore-RWKV RWKV 27.9140 37.4585 37.3314 0.9474 0.0852
5.2 CT Image Denoising Results
For CT image denoising, we compare our proposed Restore-RWKV with three specific CT image denoising methods: REDCNN [8], Eformer [9], and CTformer [10], as well as the four general image restoration methods: SwinIR [28], Restormer [43], Xformer [44], and MambaIR [29]. Table 2 indicates that Restore-RWKV achieves better CT image denoising performance with good computational efficiency compared to all other methods. Compared to the second-best method MambaIR, it achieves a 0.07 dB increase in PSNR while maintaining comparable computational complexity. Figure 5 (a) shows that the image recovered by Restore-RWKV is sharper and closer to the corresponding HQ image.
5.3 PET Image Synthesis Results
For PET image synthesis, we compare Restore-RWKV with three specific PET image synthesis methods: Xiang’s [11], DCNN [12], and ARGAN [13], as well as the four general image restoration methods: SwinIR [28], Restormer [43], Xformer [44], and MambaIR [29]. Table 3 indicates that Restore-RWKV achieves state-of-the-art performance in PET image synthesis. Compared to MambaIR, which achieves the second-best performance in PSNR and RMSE, and Xformer, which achieves the second-best results in SSIM, Restore-RWKV consistently outperforms them in all three evaluation metrics. Figure 5 (a) presents the synthesis results, showing that Restore-RWKV generates more pleasing images with clearer contrast.
@@ 5.4 All-in-One Medical Image Restoration Results
The aforementioned experimental results indicate that Restore-RWKV achieves superior performance in individual MedIR tasks. To further evaluate the model’s capacity and generalizability, we assess its performance on the all-in-one medical image restoration task, which involves using a single model to handle multiple MedIR tasks. Given the apparent disparities among different MedIR tasks, this all-in-one task is particularly challenging and serves as a robust test of model capacity and generalizability. We compare our proposed Restore-RWKV with three all-in-one image restoration methods: AirNet [16], DRMC [17], and AMIR [18], as well as four general image restoration methods: SwinIR [28], Restormer [43], Xformer [44], and MambaIR [29]. As shown in Table 4, all-in-one methods generally outperform general image restoration methods due to their specifically designed modules that address task differences. Although our Restore-RWKV does not incorporate any specialized modules for handling task disparities, it achieves the second-best results on average, slightly behind the best method, AMIR. Notably, Restore-RWKV even outperforms AMIR in PET image synthesis. These results indicate that Restore-RWKV has a strong capacity and generalizability, making it a robust backbone for medical image restoration.
[…]
Method Type Params (M) FLOPs (G) MRI Image Super-Resolution CT Image Denoising PET Image Synthesis Average
PSNR↑ SSIM↑ RMSE↓ PSNR↑ SSIM↑ RMSE↓ PSNR↑ SSIM↑ RMSE↓ PSNR↑ SSIM↑ RMSE↓
[…]
DRMC Transformer 0.6236 9.9154 29.5466 0.9032 38.1691 33.2770 0.9153 8.8674 36.1909 0.9376 0.0960 33.0048 0.9187 15.7108
[…]
SwinIR Transformer 11.4977 187.9274 30.7966 0.9235 33.1968 33.5414 0.9162 8.5991 37.0070 0.9437 0.0880 33.7817 0.9278 13.9613
[…]
MambaIR Mamba 31.5035 34.3491 31.3145 0.9305 31.3150 33.5025 0.9165 8.6345 37.1669 0.9458 0.0864 33.9946 0.9309 13.3453
[…]
Figure 5: Visual comparison of different methods. (a) MRI image super-resolution. (b) CT image denoising. (c) PET image synthesis. Zoomed ROI of the rectangle region is recommended for better visualization. Arrows in (b) indicate regions with notable differences.
@@ 5.5 Ablation Studies
We conduct ablation experiments on the MRI image super-resolution task to investigate the significance of two key innovations in the Restore-RWKV: Re-WKV attention and Omni-Shift mechanisms.
Effect of Re-WKV and Omni-Shift. To evaluate the effectiveness of Re-WKV attention and Omni-Shift, we conduct experiments by replacing them with other WKV attention layers, such as Uni-WKV [30] and Bi-WKV [32], as well as token shift layers, including Uni-Shift [30] and Quad-Shift [32]. Table 5 indicates that both Re-WKV and Omni-Shift can effectively improve model performance, with the combination of both achieving the best results. Notably, compared to the original RWKV’s [30] combination of ”Uni-WKV + Uni-Shift,” our proposed combination ”Re-WKV + Omni-Shift” achieves an improvement of over 0.77 dB in PSNR. Figure 6 illustrates the impact of different attention and token shift combinations on the effective receptive field (ERF) of models. Generally, a larger receptive field allows the model to capture information from a wider region, thereby enhancing restoration performance. We can see that the original RWKV’s [30] combination of ”Uni-WKV + Uni-Shift” achieves only a local receptive field, while the Vision-RWKV’s [32] combination of ”Bi-WKV + Quad-Shift” achieves a global receptive field. Our proposed Restore-RWKV’s combination of ”Re-WKV + Omni-Shift” achieves the largest global receptive field. This indicates that the combination of ”Re-WKV + Omni-Shift” most effectively ensures the model’s ability to capture global dependencies in 2D images.
Table 5: Component analysis of Re-WKV and Omni-Shift. The best results are highlighted in bold.
Component PSNR↑ SSIM↑ RMSE↓
Uni-WKV + Uni-Shift 31.3184 0.9315 31.3278
Bi-WKV + Q-Shift 31.7261 0.9359 30.1125
Re-WKV + Q-Shift 31.9273 0.9386 29.4302
Bi-WKV + Omni-Shift 31.9377 0.9384 29.4318
Re-WKV + Omni-Shift (Ours) 32.0913 0.9408 28.9713
@@
Table 6: Ablation study on the recurrence number in Re-WKV. The best results are highlighted in bold.
@@
$M=1$ (H-Scan) 31.9377 0.9384 29.4318
$M=1$ (V-Scan) 31.9337 0.9383 29.4371
$M=2$ (Ours) 32.0913 0.9408 28.9713
$M=3$ 32.1082 0.9409 28.8183
$M=4$ 32.1133 0.9409 28.7891
Table 7: Ablation study on architecture designs for conducting attention from both horizontal and vertical scan directions. The best results are highlighted in bold.
@@
Alternating between Blocks 31.9728 0.9393 29.2942
Sum in a Block 31.8968 0.9386 29.4702
Recurrence in a Block (Ours) 32.0913 0.9408 28.9713
Table 8: Ablation study on Omni-Shift. The best results are highlighted in bold.
@@ 6 Conclusion
In this paper, we propose Restore-RWKV, leveraging the advanced RWKV architecture for the first time in the field of medical image restoration (MedIR). To achieve superior restoration performance, Restore-RWKV innovatively incorporates a recurrent WKV (Re-WKV) attention mechanism with linear computational complexity and an omnidirectional token shift (Omni-Shift) mechanism to effectively capture global and local dependencies in 2D images, respectively. Extensive experiments on various MedIR tasks demonstrate that our proposed Restore-RWKV is an efficient and effective backbone for medical image restoration.

### v3

@@ 4.2 Implementation
For the architecture of Restore-RWKV, the number of R-RWKV blocks is set to $N_{1}=N_{\text{refinement}}=4$, $N_{2}=N_{3}=6$, and $N_{4}=8$. The number of input channels is $C=48$, and the number of attention recurrences in Re-WKV is $M=2$. We also introduce two additional model variants: the lightweight variant, Restore-RWKV-light, and the all-in-one medical image restoration variant, Restore-RWKV-routing. The Restore-RWKV-light differs from the standard Restore-RWKV in both the number of blocks and channels, with values set to $N_{1}=N_{2}=N_{4}=N_{\text{refinement}}=1$, $N_{3}=4$ and $C=16$. The Restore-RWKV-routing variant adopts the task-adaptive routing strategy introduced by the AMIR model [22], integrating a spatial routing module at the beginning of each encoder and decoder level in Restore-RWKV. To maintain parameter efficiency, the channel number is reduced to $C=40$
@@ 4.3 Evaluation
To evaluate image quality, we use three widely recognized metrics for quantitative comparisons: Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM), and Root Mean Squared Error (RMSE). Higher PSNR and SSIM values indicate better performance, while lower RMSE values signify improved results. To further assess model performance on clinically relevant regions of interest (ROIs), a radiologist annotates the liver and lesion in the testing dataset of PolarStar m660 for PET image synthesis. PSNR is then computed specifically over these ROIs. The liver is selected because it is commonly used to evaluate the uniformity and noise levels of PET images, while the lesion is prioritized as it represents the most critical region for disease diagnosis. To evaluate model efficiency, we calculate the number of model parameters and floating-point operations per second (FLOPs). The input image size is set to 128$\times$128 when calculating FLOPs.
@@ 5 Experimental Results
5.1 PET Image Synthesis Results
For PET image synthesis, we compare Restore-RWKV with seven PET image synthesis methods: Xiang’s model [1], DCNN [3], Xu’s model [2], CycleWGAN [4], DCITN [5], ARGAN [6] and DRMC [8]. Table 1 indicates that Restore-RWKV significantly (p-value $<$ 0.05 in the t-test) outperforms other methods on both the in-distribution PolarStar m660 dataset and out-of-distribution PolarStar Flight dataset. This demonstrates the effectiveness and strong model generalization ability of the proposed Restore-RWKV. Notably, even the lightweight variant Restore-RWKV-light can already outperform those comparison methods. Fig. 5 presents the visual comparison results. It is clear that the comparison methods generally fail to differentiate between noise and image details, resulting in the smoothing of important features during denoising. In contrast, the proposed Restore-RWKV is more effective at preserving these details, such as the sharpness of spine contours and the contrast in high-uptake regions. This advantage stems from Restore-RWKV’s global receptive field, which enables it to leverage both local and global contexts to better distinguish between noise and image details, leading to superior recovery of fine details.
To further evaluate the model’s effectiveness and generalization ability, we assess its performance on clinical ROIs and low-dose PET images with varying DRFs. Figure 6 (a) shows the model’s performance on clinical ROIs of the liver and lesion. The proposed Restore-RWKV method most effectively preserves these two regions, especially the lesion region, where it outperforms all comparison methods by a large margin. Figure 6 (b) illustrates the model’s performance on low-dose PET images with different DRFs. It is evident that Restore-RWKV consistently performs the best across all DRF levels, demonstrating its strong generalization ability.
Table 2: CT image denoising results. The best result is marked in bold, and the second-best result is underlined.
AAPM LDCT
Method Params (M) FLOPs (G) PSNR↑ SSIM↑ RMSE↓ PSNR↑ SSIM↑ RMSE↓
REDCNN 1.85 24.92 33.19 0.9113 8.9427 35.07 0.9201 7.2062
WGAN-VGG 0.06 0.92 31.96 0.8977 10.3366 33.72 0.9004 8.4734
EDCNN 0.08 1.32 33.41 0.9155 8.7401 35.37 0.9242 6.9923
TransCT 13.23 0.65 32.62 0.9082 9.5330 34.64 0.9183 7.5752
Eformer 0.34 1.28 33.35 0.9175 8.8030 35.48 0.9262 6.9151
CTformer 1.45 3.44 33.25 0.9134 8.8974 35.31 0.9240 7.0475
DenoMamba 112.62 110.24 33.53 0.9149 8.6115 35.67 0.9259 6.7565
@@ 5.1 PET Image Synthesis Results
Restore-RWKV 27.91 37.46 33.80 0.9198 8.3600 36.06 0.9310 6.4614
Figure 7: Visual comparison of different methods in CT image denoising. The zoomed-in rectangular region is recommended for a better comparison.
5.2 CT Image Denoising Results
For CT image denoising, we compare our proposed Restore-RWKV with seven CT image denoising methods: REDCNN [9], WGAN-VGG [10], EDCNN [11], TransCT [12], Eformer [13], CTformer [14], and DenoMamba [15]. Table 2 indicates that Restore-RWKV achieves the best CT image denoising performance (p-value $<$ 0.05 in the t-test) compared to all other methods on both the in-distribution dataset AAPM and the out-of-distribution dataset LDCT. Even the Restore-RWKV-light, with only 1.16 million parameters, outperforms the best-performing comparison method, DenoMamba, which has 112.62 million parameters. These results indicate that Restore-RWKV is an efficient and effective model for CT denoising. The visual comparison result in Fig. 7 (a) shows that the image recovered by Restore-RWKV is sharper and clearer.
Table 3: MRI image super-resolution results. The best result is marked in bold, and the second-best result is underlined.
IXI HCP
Method Params (M) FLOPs (G) PSNR↑ SSIM↑ RMSE↓ PSNR↑ SSIM↑ RMSE↓
VDSR 0.66 10.89 30.04 0.9140 36.0508 27.74 0.8191 77.1382
DAGAN 98.59 33.98 30.55 0.9189 34.0866 28.90 0.8630 67.5468
FAWDN 5.07 82.92 30.04 0.9136 35.9564 28.97 0.8665 67.0522
SwinMR 11.40 200.18 30.93 0.9253 32.7339 29.03 0.8666 66.5876
SDAUT 67.23 55.25 30.96 0.9257 32.5928 29.08 0.8710 66.2161
F-UNet 32.12 10.21 31.26 0.9314 31.5675 29.08 0.8697 66.1355
@@ IXI HCP
Restore-RWKV-light 1.16 1.52 31.36 0.9309 31.2564 29.22 0.8738 65.1056
Restore-RWKV 27.91 37.46 32.09 0.9408 28.9713 29.41 0.8788 63.8063
Figure 8: Visual comparison of different methods in MRI image super-resolution. The zoomed-in rectangular region is recommended for a better comparison
Table 4: All-in-one medical image restoration results. The best result is marked in bold, and the second-best result is underlined.
PET Image Synthesis CT Image Denoising MRI Image Super-Resolution
Method Params (M) FLOPs (G) PSNR↑ SSIM↑ RMSE↓ PSNR↑ SSIM↑ RMSE↓ PSNR↑ SSIM↑ RMSE↓
ARGAN 31.14 8.48 36.75 0.9389 0.0907 32.92 0.9111 9.2110 30.08 0.9083 35.7999
DenoMamba 112.62 110.24 36.81 0.9367 0.0895 33.18 0.9115 8.9512 30.32 0.9091 34.6972
@@
[…]
DRMC 0.62 9.92 36.19 0.9376 0.0960 33.28 0.9153 8.8674 29.55 0.9032 38.1691
[…]
Restore-RWKV-routing 24.85 31.95 37.32 0.9479 0.0852 33.74 0.9189 8.4148 32.10 0.9408 28.9507
5.3 MRI Image Super-Resolution Results
For MRI image super-resolution, we compare our proposed Restore-RWKV with seven image super-resolution methods: VDSR [33], DAGAN [16], FAWDN [17], SwinMR [18], SDAUT [19], F-Unet [20], and MambaIR [32]. Table 3 demonstrates that Restore-RWKV significantly (p-value $<$ 0.05 in the t-test) outperforms all comparison methods on both the in-distribution dataset IXI and the out-of-distribution dataset HCP in MRI image super-resolution. The Mamba-based model, MambaIR, which is also reported to achieve a global receptive field with linear complexity, achieves the second-best result. However, as illustrated in Fig.1, Restore-RWKV can achieve a more significant global receptive field than MambaIR and, as a result, better captures global dependencies to recover image structures and details. Fig.8 presents visualization results across different methods, highlighting that Restore-RWKV excels in recovering realistic MRI image details.
@@ 5.4 All-in-One Medical Image Restoration Results
The aforementioned experimental results indicate that Restore-RWKV achieves superior performance in individual MedIR tasks. To further evaluate the model’s capacity and generalization ability, we assess its performance on the all-in-one medical image restoration task, which involves using a single model to handle multiple MedIR tasks. Given the apparent disparities among different MedIR tasks, this all-in-one task is particularly challenging and serves as a robust test of model capacity and generalization ability. For a fair comparison with all-in-one models, we also introduce a model variant Restore-RWKV-routing, which incorporates the task-adaptive routing strategy developed by the AMIR model [22]. We compare our proposed Restore-RWKV and Restore-RWKV-routing with three best-performing comparison methods in each individual task: ARGAN [6], DenoMamba [15], and MambaIR [32], as well as three all-in-one image restoration methods: AirNet [21], DRMC [8], and AMIR [22]. As shown in Table 4, Restore-RWKV outperforms the three best-performing comparison methods from individual tasks, indicating that it has the best model capacity and generalization ability to deal with different tasks. Meanwhile, Restore-RWKV-routing achieves the best overall results, surpassing all other all-in-one methods. This suggests that Restore-RWKV is a robust restoration backbone that can be easily extended to multi-task settings through the integration of specific modules.
@@ 5.5 Ablation Studies
We conduct ablation experiments on the IXI dataset for MRI image super-resolution task to investigate the significance of two key innovations in the Restore-RWKV: Re-WKV attention and Omni-Shift mechanisms.
Effect of Re-WKV and Omni-Shift. To evaluate the effectiveness of Re-WKV attention and Omni-Shift, we conduct experiments by replacing them with other WKV attention layers, such as Uni-WKV [35] and Bi-WKV [37], as well as token shift layers, including Uni-Shift [35] and Quad-Shift [37]. Table 5 indicates that both Re-WKV and Omni-Shift can effectively improve model performance, with the combination of both achieving the best results. Notably, compared to the original RWKV’s [35] combination of ”Uni-WKV + Uni-Shift,” our proposed combination ”Re-WKV + Omni-Shift” achieves an improvement of over 0.77 dB in PSNR. Fig. 9 illustrates the impact of different attention and token shift combinations on the effective receptive field (ERF) of models. Generally, a larger receptive field allows the model to capture information from a wider region, thereby enhancing restoration performance. We can see that the original RWKV’s [35] combination of ”Uni-WKV + Uni-Shift” achieves only a local receptive field, while the Vision-RWKV’s [37] combination of ”Bi-WKV + Quad-Shift” achieves a global receptive field. Our proposed Restore-RWKV’s combination of ”Re-WKV + Omni-Shift” achieves the largest global receptive field. This indicates that the combination of ”Re-WKV + Omni-Shift” most effectively ensures the model’s ability to capture global dependencies in 2D images.
@@
Table 5: Component analysis of Re-WKV and Omni-Shift. The best results are highlighted.
Method PSNR↑ SSIM↑ RMSE↓
Uni-WKV + Uni-Shift 31.32 0.9315 31.3278
Bi-WKV + Q-Shift 31.73 0.9359 30.1125
Re-WKV + Q-Shift 31.93 0.9386 29.4302
Bi-WKV + Omni-Shift 31.94 0.9384 29.4318
Re-WKV + Omni-Shift (Ours) 32.09 0.9408 28.9713
Table 6: Ablation study on the recurrence number in Re-WKV. The best results are highlighted.
@@
$M=1$ (H-Scan) 31.94 0.9384 29.4318
$M=1$ (V-Scan) 31.93 0.9383 29.4371
$M=2$ (Ours) 32.09 0.9408 28.9713
$M=3$ 32.11 0.9409 28.8183
$M=4$ 32.11 0.9409 28.7891
Table 7: Ablation study on architecture designs for conducting attention from both horizontal and vertical scan directions. The best results are highlighted.
@@
Alternating between Blocks 31.97 0.9393 29.2942
Sum in a Block 31.90 0.9386 29.4702
Recurrence in a Block (Ours) 32.09 0.9408 28.9713
Table 8: Ablation study on Omni-Shift. The best results are highlighted.
@@ 6 Conclusion
In this paper, we propose Restore-RWKV, leveraging the advanced RWKV architecture for the first time in the field of medical image restoration (MedIR). To achieve superior restoration performance, Restore-RWKV innovatively incorporates a recurrent WKV (Re-WKV) attention mechanism with linear computational complexity and an omnidirectional token shift (Omni-Shift) mechanism to effectively capture global and local dependencies in 2D images, respectively. These innovations make Restore-RWKV an efficient and effective backbone for medical image restoration. Notably, even a lightweight variant with only 1.16 million parameters outperforms or achieves comparable results to existing state-of-the-art methods. Extensive experiments across a range of MedIR tasks demonstrate the efficiency and effectiveness of Restore-RWKV, establishing it as a promising solution for various medical image restoration challenges. In the future, we will explore the potential of Restore-RWKV for additional MedIR tasks across other modalities and further investigate techniques to enhance both model performance and computational efficiency.
Declaration of Competing Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgments
This work is supported by the National Natural Science Foundation in China under Grant U23B2063 and 62371016, the Bejing Natural Science Foundation Haidian District Joint Fund in China under Grant L222032, the Fundamental Research Funds for the Central University of China from the State Key Laboratory of Software Development Environment in Beihang University in China, the 111 Proiect in China under Grant B13003, the SinoUnion Healthcare Inc. under the eHealth program, and the high performance computing (HPC) resources at Beihang University.